OpenStudy (anonymous):

The lengths of the sides of a right triangle are a and b, and the hypotenuse is c. Find the area of the triangle. a = 2 sqrt 3 ft.,c = 4 ft A = sq. ft.

OpenStudy (anonymous):

@Drakus

OpenStudy (anonymous):

here, we have a right angled triangle with a, b nd c as its sides. a = base(b) = \[2\sqrt{3}\] c= hypotenus (h) = 4 ft so perpendicular(b) = \[\sqrt{h^{2}-b^{2}}\] \[\sqrt{4^{2}-(2\sqrt{3)}^{2}}\] \[\sqrt{16-12}\] \[\sqrt{4}\] = 2 ft now, Area of triangle = base*perpendicular/2 = \[2\sqrt{3}\times2\div2\] \[4\sqrt{3}\div2\] \[2\sqrt{3}\]

OpenStudy (anonymous):

is the answer correct??????

OpenStudy (anonymous):

yep thanks a lot:))

OpenStudy (anonymous):

no mention dud....